# Drillboards: Adaptive Visualization Dashboards for Dynamic Personalization of Visualization Experiences

### v1

@@ 1 Introduction
Figure 1: Drillboards are adaptive visualization dashboards consisting of a hierarchy of coordinated charts. A drillboard has two modes: authoring (A) and reading (B). The hierarchy is designed in authoring mode, and then explored in reading mode. The author-designed hierarchical structure is shown in the hierarchy browser. In the reading mode, the user can drill down a chart to expand it into its constituent child charts. The user can also roll up the hierarchy to collapse several charts. Here is how the author mode (A) works: ① The author needs to select charts of their interest. ④ After selecting the operation to merge the charts and adding descriptions and title, ⑤ the chart is merged. The treeview also is updated based on the changes made (②, ③). Here is how the reader mode (B) works: ① The reader first sees a chart of their interest, and left-clicks the chart. Then, ⑥ the original chart disappears, and 3 children charts appear instead, in a highlighted manner. ⑤ By right-clicking any children charts, the charts roll-up to transform into the parent chart. Here too, the treeview (①, ③) is updated based on the changes made in the chart view.
@@
  • The DrillVis system that allows authoring of aggregate drillboards (author mode) and exploration of data via drill-down and roll-up interactions (reader mode);
@@ 2.1 Multiple Views
We draw on many of these ideas for our drillboards technique, but do not claim any contribution to the multiple view area.
@@ 2.3 Visualizing Hierarchical Aggregation
A wealth of research has been dedicated to effectively visualizing and summarizing data into hierarchical aggregates, driven primarily by the need for managing large-scale datasets efficiently. This pursuit began with the advent of data cubes [58, 36, 44, 63], which introduced algorithms aimed at optimizing computational resource utilization—notably memory and CPU usage—when handling extensive datasets. Additionally, effective summarization techniques for large-scale datasets have been explored [12]. Another focus in visualization research is on presenting data structure and relationships effectively, exemplified by methods such as trees and treemaps [55, 32, 5, 6]. Similarly, thematic representations, such as streamgraph-based visualizations [7, 18, 22], have been employed to illustrate topic evolutions [64, 11, 23].
@@ 2.4 Authoring Dashboards and Multiple Views
Because of their prevalence in business intelligence and commercial data analysis, dashboard design has long been a focus in the business community [16]. As a result, several tools support advanced authoring functionality. Tableau (née Polaris [57]) offers a powerful dashboard authoring interface integrated with its visualization and analytics capabilities. Microsoft Power BI [^1] https://www.microsoft.com/en-us/power-platform/products/power-bi is another popular choice for dashboard authoring that integrates seamlessly with other Microsoft tools and offers robust data analytics and visualization features. Launched in 2016, Google Data Studio was a free tool for creating interactive dashboards using data from various sources, including Google Analytics, Google Ads, etc. In 2022, the software suite was rebranded as Looker Studio [^2] https://lookerstudio.google.com/ after Google’s successful \$2.6B acquisition of data analytics company Looker in 2019. Finally, Qlik Sense [^3] https://www.qlik.com/us/products/qlik-sense is a data analytics platform that enables users to create interactive dashboards and reports from visualizations created during data analysis.
@@
The most recent work in this space leverages automated methods and AI to aid dashboard design. StoryFacets [46] use automatically generated dashboards as one of three storytelling mechanisms to report on data analysis outcomes. LADV [37] offers a rapid conceptualization approach for constructing dashboard templates from exemplars using a deep learning-based model, enhancing efficiency in dashboard visualization authoring for data-intensive applications. Similarly, MultiVision [65] provides a deep learning-based method that automates the design of analytical dashboards by recommending meaningful combinations of data columns and multiple charts to support data selection and composition. MEDLEY [45] automates dashboard composition by recommending collections of visualizations and filtering widgets based on specific analytical intents, simplifying dashboard authoring process. BOLT [56] automates dashboard authoring through natural language input, enabling users to express their objectives and receive relevant dashboard recommendations to streamline dashboard creation.
@@
Drillboards draw on several of these ideas, but provide a unique approach to hierarchically grouping charts into aggregates to yield an adaptive user interface. While automated methods may be utilized to customize a dashboard for a particular user, such as in QualDash [15], a drillboard at its core remains invariant—it just offers more or less detail depending on the needs of the user. This promotes consistency across users and expertise levels, and also facilitates seamless upskilling where users grow in expertise and accordingly drill deeper for more detail (as well as downskilling, in case of a user not having visited a dashboard for a longer period of time).
@@ 3.1 Design Rationale
Visualization dashboards are forms of communication-minded visualizations [61]; they are designed for use by multiple users [53]. Consider a typical visualization dashboard created for internal use in an organization or published on a public website. While the former typically has a more controlled user audience than the latter, it still remains that the multiple users that frequent a visualization will have varying expertise. This leads us to our first design goal (DG1):
@@
Third, the point of an adaptive user interface—beyond adapting to the user’s context, task, and needs—is to support users upskilling (or downskilling) using the same interface. An ideal visualization dashboard would support users learning its use over time:
@@
Finally, authoring such a one-size-fits-all adaptive dashboard must be relatively straightforward:
@@ 3.2 Model
The current drillboard view is a sequence of chart groups and atoms representing the current aggregation state of the drillboard, and represent the charts that are currently visible on screen. End-users can change the view by drilling-down into a chart group—which entails replacing the chart group with its children in the view—or rolling-up a chart atom or group—which entails replacing the chart and its siblings with their common ancestor. Many drillboards also come with pre-defined views; these are views that the drillboard author have created in advance, for example, to capture a specific level of detail or skill level.
Authoring a drillboard entails starting with the chart sequence for a regular dashboard and then building the aggregation hierarchy by iteratively applying aggregation operation that replace two or more charts (atoms or groups) with a single chart group. Figure 3 shows an example. Aggregation operations are described below.
The root of the aggregation hierarchy (the single node remaining in the yellow field in Figure 3e) and the original visualization dashboard (i.e., all of the chart atoms in the dashboard, see Figure 3a) are automatically pre-defined views. The former could theoretically be used for a complete novice—just a single chart representing all the data in the dashboard—and the latter for an expert—all data visible and no informatizon loss. In addition to building the aggregation hierarchy, authors may also define current and pre-defined views. If there is no current view, then the root of the aggregation hierarchy will form the initial view. Additional pre-defined views can also be associated with the drillboard with a specific label.
@@
Figure 3: Hierarchical aggregation. Example of how an aggregation hierarchy is formed through a sequence of aggregation operations until a single chart group remains as the root.
@@ 3.3 Aggregation Operations
All aggregation operations are essentially Merge operations that replace one or more child charts—that could either be chart groups or chart atoms—with a single chart group containing the selected children. In aggregating charts, the new parent chart group must be assigned a visual representation. The choice of group representation is typically left in the hand of the drillboard designer (DG2). There is typically—but not always—information loss inherent in this kind of merge. Here we outline several possible merge operations (Figure 2):
@@
    Label: Replace all of the chart children with a single label or text that summarizes their contents. This text could be a representative scalar, such as the mean, median, minimum, or maximum number in the underlying data. For example, several charts showing the average temperature in different regions could be replaced by the average temperature for the entire country.
@@
    Summarization: Charts representing data of similar type can be summarized using data abstraction, such as calculating an average, sum, or difference between series, yielding a single new series.
@@
    Archetype: Select one of the chart children to serve as a representative for them all. There is significant information loss in this kind of merge, but at least one of the original chart remains.
@@
    Projection: Two or more data dimensions from child charts can be projected on the axis of a scatterplot (if two) or parallel coordinate plot (if more than two). Data can be mapped to a color scale—nominal or categorical to a categorical one, and an ordinal or quantitative one to a discrete or continuous ordered scale.
@@
    Juxtaposition: Combine the child charts inside the extents of the parent chart group as small multiples [28]. This preserves all of the information in the original charts, but may lead to increased visual clutter and reduced data resolution due to reduced space.
@@
    Overlay: Overlay the child charts on top of each other [28]. This operation requires that the child charts are of the same type; e.g., two line charts with a common vertical and horizontal axes.
@@ 4 System: Drillboards
We implement the drillboards technique in a web-based interactive visualization tool called DrillVis. The tool has a basic visual interface consisting of drillboard view showing the current representation of the drillboard, and a tree view for seeing and navigating the aggregation hierarchy. The tool has two main interaction modes: (1) the author mode (see Figure 5a), and (2) the reader mode (Figure 5b). Here we describe the interface, the aggregation operators implemented, and the author and reader modes.
@@ 4.1 Visual Interface
Several options for managing the drillboard space are possible. One option is for cards to be a fixed size, which means that the grid is gradually filled out as the aggregation hierarchy is drilled down, and can eventually be scrolled to see all the charts. Another option is to use a space-filling layout, where the visualization cards are dynamically resized to always fill the screen. This may lead to a situation where the cards become too small to view properly; in this situation, automatically rolling up distant cards into groups until the size is sufficiently large (similar to SpaceTree [48]) is an option. Finally, our current layout algorithm requires that all visualization cards have the same height, but does allow for resizing individual card widths.
@@
(a) Author mode. ① is the treeview which shows hierarchical relationships between charts, and the title of charts. ② is where users can select the charts of their interest. When multiple charts are selected from the operator implementations ③, authors can choose an operation, view the result, define a new title, and write explanations.
@@ 4.2 Aggregation Operators
Select merges several charts and lets the author choose a single one among them to represent them as an archetype (Figure 7).
@@ 4.3 Author Mode
Generating charts. DrillVis supports the ingestion of tabular data in formats such as CSV or TSV, accommodating both continuous and discrete data types. For continuous data, such as time-series, the system automatically assigns time data to the $x$-axis and corresponding values to the $y$-axis, defaulting to a line chart presentation with an option to switch to a histogram. Discrete data, representing categorical information, is visualized with categories on the x-axis and their values on the y-axis in the form of bar charts. Authors can choose the type of chart between line and bar charts, or could opt for other charts. For example, authors can use scatterplots to understand the relationship between two features, as shown in Fig. 8.
We assume that the dataset contains features in the first row and time tags in the first column. However, features themselves could belong to a group (e.g., cats and dogs are part of mammals), and a single row may be insufficient to hold all the dataset metadata. Even in such scenarios, the system automatically processes these relationships.
@@
Aggregating charts. The aggregation process begins with the selection of charts to be merged (Figure 1(A) ①), aiming to create a hierarchical structure of visualizations. Consistent with Section 3.3, DrillVis provides several merge operator implementations (Figure 1(A) ④) for chart aggregation; see above. When aggregating charts, there are several rules to follow. First, charts from different datasets can only be combined if they have the same x-axis ranges or, for bar charts, the same x-axis values. Second, the system currently only supports tabular datasets and cannot yet visualize complex network graphs or advanced bespoke visualizations. We leave this as future work.
@@ 4.5 Implementation Notes
We implemented Drillboards in TypeScript using React [^4] https://react.dev/ and Node.js. Most of the user interfaces are built and designed manually by the authors using the Emotion library. [^5] https://emotion.sh/ We also further customized several components (i.e. input, menu) using MUI. [^6] https://mui.com/ The charts are generated using Recharts. [^7] https://recharts.org/ We used Cloudflare R2 for cloud storage.
@@ 4.6 Usecase Example
Bob, an expert in agriculture, wants to monitor water sensors on a farm located in Maryland, USA. He receives a time-series dataset about the farm he is located, and needs to instruct incoming interns about how to monitor water sensor data effectively. To ensure efficiency, he uses DrillVis to educate interns on key features to analyze in the water sensor data. To that end, he first builds a version for novice users to educate and instruct interns. In doing so, he uses the ‘select’ operator to indicate the important feature, and group those that exhibit similar patterns. He also creates an ‘expert’ mode, a version that contains more charts, with less abstraction. An illustration of ‘select’ in DrillVis visualization is shown in Figure 11.

### v3

@@ 1 Introduction
Figure 1: Drillboards are adaptive visualization dashboards consisting of a hierarchy of coordinated charts. A drillboard has two modes: authoring (A) and reading (B). The hierarchy is designed in authoring mode, and then explored in reading mode. The author-designed hierarchical structure is shown in the hierarchy browser. In the reading mode, the user can drill down a chart to expand it into its constituent child charts. The user can also roll up the hierarchy to collapse a few charts. Here is how the author mode (A) works: ① The author needs to select charts of their interest. ④ After selecting the operation to merge the charts and adding descriptions and title, ⑤ the chart is merged. The treeview also is updated based on the changes made (②, ③). Here is how the reader mode (B) works: ① The reader first sees a chart of their interest, and left-clicks the chart. Then, ⑥ the original chart disappears, and 3 children charts appear instead, in a highlighted manner. ⑤ By right-clicking any children charts, the charts roll-up to transform into the parent chart. Here too, the treeview (①, ③) is updated based on the changes made in the chart view.
(a) Author mode. ① is the treeview which shows hierarchical relationships between charts, and the title of charts. ② is where users can select the charts of their interest. When multiple charts are selected from the operator implementations ③, authors can choose an operation, view the result, define a new title, and write explanations. Users can view the x- and y-labels by toggling on in the button in ④.
@@
  • The DrillVis system that allows authoring of aggregate drillboards (author mode) and exploration of data via drill-down and roll-up interactions (reader mode); and
@@ 2.1 Multiple Views
We draw on many of these ideas for our drillboards technique, but do not claim any contribution to multiple views.
@@ 2.3 Visualizing Hierarchical Aggregation
A wealth of research has been dedicated to effectively visualizing and summarizing data into hierarchical aggregates, driven primarily by the need for managing large-scale datasets efficiently. This pursuit began with the advent of data cubes [59, 36, 45, 64], which introduced algorithms aimed at optimizing computational resource utilization—notably memory and CPU usage—when handling extensive datasets. Additionally, effective summarization techniques for large-scale datasets have been explored [12, 41]. Another focus is on presenting data structure and relationships effectively, exemplified by methods such as trees and treemaps [56, 32, 5, 6]. Similarly, thematic representations, such as streamgraph-based visualizations [7, 18, 22], have been employed to illustrate topic evolutions [65, 11, 23].
@@ 2.4 Authoring Dashboards and Multiple Views
Because of their prevalence in business intelligence and commercial data analysis, dashboard design has long been a focus in the business community [16]. As a result, we identify tools that support advanced authoring functionality. Tableau (née Polaris [58]) offers a powerful dashboard authoring interface integrated with its visualization and analytics capabilities. Microsoft Power BI [^1] https://www.microsoft.com/en-us/power-platform/products/power-bi is another popular choice for dashboard authoring that integrates seamlessly with other Microsoft tools and offers robust data analytics and visualization features. Launched in 2016, Google Data Studio was a free tool for creating interactive dashboards using data from various sources, including Google Analytics, Google Ads, etc. In 2022, the software suite was rebranded as Looker Studio [^2] https://lookerstudio.google.com/ after Google’s successful \$2.6B acquisition of data analytics company Looker in 2019. Finally, the Qlik Sense [^3] https://www.qlik.com/us/products/qlik-sense data analytics platform enables users to create interactive dashboards and reports from visualizations created during data analysis.
@@
The most recent work in this space leverages automated methods and AI to aid dashboard design. StoryFacets [47] use automatically generated dashboards as one of three storytelling mechanisms to report on data analysis outcomes. LADV [37] offers a rapid conceptualization approach for constructing dashboard templates from exemplars using a deep learning-based model, enhancing efficiency in dashboard visualization authoring for data-intensive applications. Similarly, MultiVision [66] provides a deep learning-based method that automates the design of analytical dashboards by recommending meaningful combinations of data columns and multiple charts to support data selection and composition. MEDLEY [46] automates dashboard composition by recommending collections of visualizations and filtering widgets based on specific analytical intents, simplifying dashboard authoring process. BOLT [57] automates dashboard authoring through natural language, enabling users to express their objectives and receive relevant recommendations to streamline dashboard creation.
@@
Drillboards provides a unique approach to hierarchically grouping charts into aggregates to yield an adaptive user interface. While automated methods may be utilized to customize a dashboard for a particular user, such as in QualDash [15], a drillboard at its core remains invariant—it just offers more or less detail depending on the needs of the user. This promotes consistency across users and expertise levels, and also facilitates seamless upskilling where users grow in expertise and accordingly drill deeper for more detail (as well as downskilling, in case of a user not having visited a dashboard for a longer period of time).
@@ 3.1 Design Rationale
Visualization dashboards are forms of communication-minded visualizations [62]; they are designed for use by multiple users [54]. Consider a typical visualization dashboard created for use in an organization or published on a website. While the former typically has a more controlled user audience than the latter, it still remains that the multiple users that frequent a visualization will have varying expertise. This leads us to our first design goal (DG1):
@@
Third, the point of an adaptive user interface—beyond adapting to the user’s context, task, and needs—is to support users upskilling (or downskilling—when the user did not use the tool for a long time, and needs time to familiarize with the tool) using the same interface. An ideal visualization dashboard would support users learning its use over time:
@@
Finally, authoring such a one-size-fits-all adaptive dashboard must be relatively straightforward. For this, the tool must be able to effectively guide users to author, and provide a good overview of how the hierarchy is made:
@@ 3.2 Model
The current drillboard view is a sequence of piles and atoms representing the current aggregation state of the drillboard, and represent the charts that are currently visible on screen. End-users can change the view by drilling-down into a pile—which entails replacing the pile with its children in the view—or rolling-up a chart atom or group—which entails replacing the chart and its siblings with their common ancestor. Many drillboards also come with pre-defined views; these are views that the drillboard author have created in advance, for example, to capture a specific level of detail or skill level.
Authoring a drillboard entails starting with the chart sequence for a regular dashboard and then building the aggregation hierarchy by iteratively applying aggregations that replace two or more charts (atoms or groups) with a single pile. Figure 4 shows an example. Aggregation operations are described below.
The root of the aggregation hierarchy (the single node remaining in the yellow field in Figure 4e) and the original visualization dashboard (i.e., all of the chart atoms in the dashboard, see Figure 4a) are automatically pre-defined views. The former could theoretically be used for a complete novice—just a single chart representing all the data in the dashboard—and the latter for an expert—all data visible and no information loss. In addition to building the aggregation hierarchy, authors may also define current and pre-defined views (DG1). If there is no current view, then the root of the aggregation hierarchy will form the initial view. Additional pre-defined views can also be associated with the drillboard with a specific label.
@@
Figure 4: Hierarchical aggregation. Example of how an aggregation hierarchy is formed through a sequence of aggregation operations until a single pile remains as the root.
@@ 3.3 Aggregation Operations
All aggregation operations are essentially Merge operations that replace one or more child charts—that could either be piles or chart atoms—with a single pile containing the selected children. In aggregating charts, the new parent pile must be assigned a visual representation. The choice of group representation is typically left in the hand of the drillboard designer (DG2). There is typically—but not always—information loss inherent in this kind of merge. Here we outline 6 possible merge operations (Figure 3.) The list provided here consists of elements that support the contextualization of charts by integrating child charts. We mainly focus on three types of merge operations (DG2): (1) operations for structural organization and abstraction (SOA), (2) comparative and relational analysis (CRA), and (3) integrated representation (IR). This list is not exhaustive and can be extended:
@@
    Label (SOA): Replace all of the chart children with a single label or text that summarizes their contents (i.e., not a form of data visualization). This text could be a representative scalar, such as the mean, median, minimum, or maximum number in the underlying data. For example, charts showing the temperature in different regions could be replaced by the average temperature for the country.
@@
    Summarization (SOA): Charts representing data of similar type can be summarized using data abstraction, such as calculating an average, sum, or difference between series, yielding a single new series.
@@
    Archetype (SOA): Select one of the chart children to serve as a representative for them all. There is significant information loss in this kind of merge, but at least one of the original chart remains.
@@
    Projection (CRA, IR): Two or more data dimensions from child charts can be projected on the axis of a scatterplot (if two) or parallel coordinate plot (if more than two). Data can be mapped to a color scale—nominal or categorical to a categorical one, and an ordinal or quantitative one to a discrete or continuous ordered scale.
@@
    Juxtaposition (CRA): Combine the child charts inside the extents of the parent pile as small multiples [28]. This preserves all of the information in the original charts, but may lead to increased visual clutter and reduced data resolution due to reduced space.
@@
    Overlay (IR): Overlay the child charts on top of each other [28]. This operation requires that the child charts are of the same type; e.g., two line charts with a common vertical and horizontal axes.
@@ 4 System: Drillboards
We implement the drillboards technique in a web-based interactive visualization tool called DrillVis. The tool has a basic visual interface consisting of drillboard view showing the current representation of the drillboard, and a tree view for seeing and navigating the aggregation hierarchy. The tool has two main interaction modes: (1) the author mode (see Figure 2a), and (2) the reader mode (Figure 2b). Here we describe the interface, the aggregation operators, and the author and reader modes.
@@ 4.1 Visual Interface
Three options for managing the drillboard space are possible. One option is for cards to be a fixed size, which means that the grid is gradually filled out as the aggregation hierarchy is drilled down, and can eventually be scrolled to see all the charts. Another option is to use a space-filling layout, where the visualization cards are dynamically resized to always fill the screen. This may lead to a situation where the cards become too small to view properly; in this situation, automatically rolling up distant cards into groups until the size is sufficiently large (similar to SpaceTree [49]) is an option. Finally, our current layout algorithm requires that all visualization cards have the same height, but does allow for resizing individual card widths.
@@ 4.2 Aggregation Operators
Select merges multiple charts and lets the author choose a single one to represent them as an archetype (Figure 7).
@@ 4.3 Author Mode
Generating charts. DrillVis supports the ingestion of tabular data in formats such as CSV or TSV, accommodating both continuous and discrete data types. For continuous data, such as time-series, the system automatically assigns time data to the $x$-axis and corresponding values to the $y$-axis, defaulting to a line chart presentation with an option to switch to a histogram. Discrete data, representing categorical information, is visualized with categories on the x-axis and their values on the y-axis in the form of bar charts. Authors can choose the type of chart between line and bar charts, or could opt for other charts. For example, authors can use scatterplots to understand the relationship between two features, as shown in Figure 8.
We assume that the dataset contains features in the first row and time tags in the first column. However, features themselves could belong to a group (e.g., cats and dogs are part of mammals), and a single row may be insufficient to hold all the dataset metadata. Even in such scenarios, the system automatically processes these relationships (DG4).
@@
Aggregating charts. The aggregation process begins with the selection of charts to be merged (Figure 1(A) ①), aiming to create a hierarchical structure of visualizations. Consistent with Section 3.3, DrillVis provides various merge operator implementations (Figure 1(A) ④) for chart aggregation; see above. When aggregating charts, there are rules to follow. First, charts from different datasets can only be combined if they have the same x-axis ranges or, for bar charts, the same x-axis values. We automatically disable inapplicable aggregation operations by graying them out. (see Figure 2(a) ③) (DG4). Furthermore, the treeview interface is displayed at the left to facilitate checking of how the hierarchy is made (DG4). Second, the system currently only supports tabular datasets and cannot yet visualize complex network graphs or advanced bespoke visualizations. We leave this as future work.
@@
Constraints in aggregation. For many operations, certain constraints must be met. For example, calculations such as averages, sums, or differences are only possible when both the x- and y-axes of the charts share the same dimensions. For projection, the x-axis of both charts must share the same dimension. Multiple lines can be overlaid in a single chart if the two charts also share the same dimensions on the y-axis.
Even though we used relatively simple charts, we encounter numerous scenarios where the merging process is not applicable. For example, consider a chart that is already merged as a juxtaposition, alongside another chart that overlays two visualizations. In this situation, we identify a few operations that are not possible: arithmetic summarization, projection, and additional overlays. Juxtaposition may still work, but only until the unit charts become too small to be legible. Another example involves aggregating two charts, one of which is merged through text labeling, and the other through projection. While aggregation through labeling or an archetype might be feasible, many arithmetic operations or overlays are not. In our case, we have automatically disabled such merge operations.
Consistency in aggregation policies. We may face confusion when the meaning of these aggregations is ambiguous. For instance, should we allow the arithmetic merging of a leaf line chart with another line chart previously merged as an ‘archetype’ using two line charts and one labeled chart? In our case, we decided against this operation because the resulting aggregation would be difficult to interpret.
We do not claim that our approach is the only solution. Depending on the goals of Drillboards, policies may vary. We argue that consistency is crucial to avoid confusing users, so well-crafted guidelines are needed to determine when merging is permitted and when it is not.
@@ 4.5 Implementation Notes
We implemented Drillboards in TypeScript using React [^4] https://react.dev/ and Node.js. The user interface is built and designed manually by the authors using the Emotion library. [^5] https://emotion.sh/ We also further customized components (i.e. input, menu) using MUI. [^6] https://mui.com/ The charts are generated using Recharts. [^7] https://recharts.org/ We used Cloudflare R2 for cloud storage.
@@ 4.6 Usecase Example
Bob, an expert in agriculture, wants to monitor water sensors on a farm located in Maryland, USA. He receives a time-series dataset about the farm he is located, and needs to instruct incoming interns about how to monitor water sensor data effectively. To ensure efficiency, he uses DrillVis to educate interns on key features to analyze in the water sensor data. To that end, he first builds a version for novice users to educate and instruct interns. In doing so, he uses the ‘select’ operator to indicate the important feature, and group those that exhibit similar patterns. He also creates an ‘expert’ mode, a version that contains more charts, with less abstraction. An illustration of ‘select’ is shown in Figure 11.
When new interns arrive, Bob instructs them on monitoring soil conditions. He begins with key features at the top of the hierarchy using novice mode, then drills down (as shown in Figure 11) to detail soil data. Novice users start with basic information, but as they gain experience, they must be ready to review more detailed data.
After guiding novice interns, Bob instructs seasoned interns using the same drillboards hierarchy. For them, he selects expert mode to explain advanced features.